A man weighs 180lb on the earth's surface. What would he weigh at a distance from the center of the earth of 3times the radius of the earth?


1
Expert's answer
2021-03-24T19:38:52-0400
"W=G\frac{mM}{R^2}\\W'=G\frac{mM}{(3R)^2}=\frac{W}{9}
\\W'=\frac{180}{9}=20\ lb"